# Causal Fingerprints of AI Generative Models

### v1

@@ 1 Introduction
The rapid evolution of generative models has significantly improved AI-generated content (AIGC), particularly in producing highly realistic images. However, this creates challenges for model attribution, which aims to identify the correct source model that generated an image. Model attribution is crucial for AIGC safety [1]: it offers an auditing mechanism of image authenticity to counter malicious forgeries [2, 3]; meanwhile, it makes the source model traceable in its outputs, thereby safeguarding the model owner’s copyright from pirates[4, 5, 6].
@@
Prior work typically defines model fingerprints as residual noise or image artifacts introduced by the generator and adopts a feature-extraction paradigm. While effective for identifying specific models, these methods depend heavily on model-specific cues and generalize poorly to models that differ in these predefined features. For instance, some Generative Adversarial Networks (GANs) produce checkerboard artifacts that have been exploited as fingerprints in prior studies [15, 16]; however, this cue does not transfer to models that exhibit no checkerboard patterns, such as some diffusion models. We argue that such feature-based approaches yield fragmentary and limited fingerprints, while the comprehensive fingerprint that captures true causality between image provenance and model traces, what we term as the causal fingerprint, remains largely overlooked.
@@
To close this gap, we present the first definition of causal fingerprint (CF) of AI generative models, together with a causality-inspired disentanglement framework that separates CFs from image-specific content and style. The framework extracts CFs within a semantics-invariant latent space derived from pretrained diffusion reconstruction residuals [12], and further increases fingerprint granularity by exploring and embedding broader, more diverse feature representations [18]. Moreover, the causal properties of the extracted fingerprints enable the construction of counterfactual fingerprints that facilitate image source obfuscation, based on which we design a model anonymization method using fingerprint-constrained PGD adversarial perturbations. To assess the causality of the fingerprints and their utility for model attribution, we conduct experiments on a challenging AI-generated image benchmark and compare against six representative baselines from different categories for attributing four generative models, including GANs and diffusion models. Additional analyses, including an ablation study on feature representations for fine-grained fingerprints, fingerprint visualizations, and model anonymization using counterfactual fingerprints, are performed to further validate the presence of fingerprints and their causality. Experimental results demonstrate that our decoupling framework successfully extracts CFs, surpasses prior methods in model attribution, and achieves model anonymization via counterfactual examples, indicating strong potential for forgery detection, model copyright tracing, and identity protection.
@@ 2.1 Definition of Causal Fingerprint (CF)
Here, $f(A)$ denotes the mapping function that extracts fingerprints from artifact $A$, wherein artifact $A$ captures specific traces during the model training process, independent of content $C$ and style $S$. To further enrich the fingerprint representation, causal fingerprints are defined by weighting and concatenating the differences in projections across multiple semantically invariant embedding spaces, as shown in Equation 2.
where $\mathcal{E}$ denotes the set of semantically invariant embedding spaces, $\phi_{s}(\cdot)$ represents the feature extraction function for the $s$th embedding space, $w_{s}$ denotes the weighting coefficient, and $r_{dire}$ denotes the residual computed via the pre-trained Diffusion-based Residual Model (DIRE), as defined in detail in Sec.2.2.
@@ 2.2 Decouple CF in Semantic-Invariant Latent Space
Among these, $\hat{X}_{dire}$ represents the image reconstructed by the DIRE model, which is used for calculating the causal fingerprint $F_{G}$ in Sec.2.1.
@@ 2.4 Network Architecture
Here, $x\in\mathbb{R}^{N}$ represents the original 128-vector feature, while $\hat{x}\in\mathbb{R}^{N}$ denotes the fingerprint feature vector reconstructed by the variational autoencoder (VAE). $x_{i}$ and $\hat{x}_{i}$ respectively denote the $i$-th element of the original fingerprint feature and the reconstructed fingerprint feature. Additionally, $\mu\in\mathbb{R}^{K}$ denotes the mean vector in the latent space, $\sigma^{2}\in\mathbb{R}^{K}$ denotes the variance vector in the latent space, and the $KL$ divergence between the normally distributed output of the encoder $\mathcal{N}(\mu,\sigma^{2})$ and the standard normal distribution $\mathcal{N}(0,1)$.
Model Attribution and Anonymization. Our attribution network aims to predict the source model of AI-generated images. It takes the causal fingerprint of an image obtained through VAE optimization as input and predicts the identity of its source model. The attribution network we employ utilizes a pre-trained CLIP [23] (ViT-B/32 architecture) enhanced with an attention mechanism [24] to bolster classification capabilities. It is fine-tuned using standard cross-entropy loss. The architecture diagram in Fig 2 illustrates our model’s artifact extraction, causal fingerprint optimization, visual disentanglement, and attribution process.
In the counterfactual approach, we employ the Projected Gradient Descent (PGD) adversarial perturbation algorithm[25]. By applying perturbations constrained by the causal fingerprint to images, we alter model predictions while maintaining perturbations within the $\mathcal{L}^{\infty}$ norm constraint, achieving anonymization. Its significance lies in: concealing model identity to protect privacy, enhancing attribution analysis accuracy and robustness, isolating causal effects to reduce bias, and supporting secure and reliable generative model development.
@@ 3.1 Experimental Setup
Datasets. To evaluate the performance of model attribution using fingerprints across diverse environments, we constructed the GM-GenImage dataset. Based on GenImage[26], it comprises AI-generated images trained on authentic ImageNet images, offering diverse content and strong generalisability. We selected 10K images each from four generative models—BigGAN, ProGAN, Glide, and Stable Diffusion(SD)—totalling 40K images to construct GM-GenImage. This dataset avoids the computational overhead of training multiple models independently, thereby conserving computational resources and reducing computational load for subsequent diffusion model training and causal decoupling of diffusion reconstruction residuals.
Baseline. We consider four major categories of existing generative model attribution methods: those based on pixel-level color, frequency domain features, supervised learning, and manifold learning features. We evaluate representative methods from each group and compare them with our proposed approach, including methods based on pixel-level colour: McCloskey et al. [16], and Hae et al. [18]; frequency domain feature-based methods: Dzanic[27]; supervised learning-based methods: Yu et al. [14], and Wang et al. [13]; Euclidean manifold learning-based methods: Hae et al. [28]. Their respective embedding spaces are utilized for comparative analysis.
@@ 3.2 Results
Table 1 presents the model attribution results for fingerprint recognition methods. The convolutional neural network (CNN) classifier outperforms KNN and support vector machines (SVM). Ablation experiments indicate that both the pre-trained ViT model’s class-labeled features and the grayscale FFT low-frequency spatial extraction features lead to poor fingerprint attribution performance. Subsequent multi-space fusion techniques improved attribution performance by reducing weights. Additionally, the FDR metric validated the separability of fingerprint feature spaces, with higher FDR values indicating stronger attribution capability. This method achieves significantly better FDR values than competitors, with an average attribution accuracy improvement of 22.78% (compared to 33.61%, 16.17%, 14%, 15.97%, 23.54%, and 33.42% for other methods), confirming the effectiveness of causal fingerprints in generative model attribution.
@@
Yu et al. (ICCV 2019) [14] 64.43 148.16 0.657 0.644
Wang et al. (CVPR 2019) [13] 81.87 3.49 0.837 0.819
McCloskey et al. (arXiv 2018) [16] 84.04 3.38 0.843 0.840
Dzanic et al. (NeurIPS 2020) [27] 82.13 48.40 0.824 0.821
Hae et al. (CVPR 2024) [18] 64.62 1.24 0.699 0.646
Hae et al. (arXiv 2025) [28] 74.5 83.34 0.858 0.745
@@
multi_fingerprint($F_{0}$) 94.33 312.77 0.944 0.943
multi_fingerprint($F_{1}$) 98.04 357.01 0.980 0.980
multi_fingerprint($F_{2}$) 88.83 199.67 0.888 0.888

### v2

@@ 1 Introduction
The rapid evolution of generative models has significantly improved AI-generated content (AIGC), particularly in producing highly realistic images. However, this creates challenges for model attribution, which aims to identify the correct source model that generated an image. Model attribution is crucial for AIGC safety [13, 14]: it offers an auditing mechanism of image authenticity to counter malicious forgeries [26, 9]; meanwhile, it makes the source model traceable in its outputs, thereby safeguarding the model owner’s copyright from pirates[6, 22, 20].
@@
Prior work typically defines model fingerprints as residual noise or image artifacts introduced by the generator and adopts a feature-extraction paradigm. While effective for identifying specific models, these methods depend heavily on model-specific cues and generalize poorly to models that differ in these predefined features. For instance, some Generative Adversarial Networks (GANs) produce checkerboard artifacts that have been exploited as fingerprints in prior studies [18, 19, 15]; however, this cue does not transfer to models that exhibit no checkerboard patterns, such as some diffusion models. We argue that such feature-based approaches yield fragmentary and limited fingerprints, while the comprehensive fingerprint that captures true causality between image provenance and model traces, what we term as the causal fingerprint, remains largely overlooked.
@@
To close this gap, we present the first definition of causal fingerprint (CF) of AI generative models, together with a causality-inspired disentanglement framework that separates CFs from image-specific content and style. The framework extracts CFs within a semantics-invariant latent space derived from pretrained diffusion reconstruction residuals [28], and further increases fingerprint granularity by exploring and embedding broader, more diverse feature representations [25, 14, 16]. Moreover, the causal properties of the extracted fingerprints enable the construction of counterfactual fingerprints that facilitate image source obfuscation, based on which we design a model anonymization method using fingerprint-constrained PGD adversarial perturbations. To assess the causality of the fingerprints and their utility for model attribution, we conduct experiments on a challenging AI-generated image benchmark and compare against six representative baselines from different categories for attributing four generative models, including GANs and diffusion models. Additional analyses, including an ablation study on feature representations for fine-grained fingerprints, fingerprint visualizations, and model anonymization using counterfactual fingerprints, are performed to further validate the presence of fingerprints and their causality. Experimental results demonstrate that our decoupling framework successfully extracts CFs, surpasses prior methods in model attribution, and achieves model anonymization via counterfactual examples, indicating strong potential for forgery detection, model copyright tracing, and identity protection.
@@ 2.1 Definition of Causal Fingerprint (CF)
Here, $f(A)$ denotes the mapping function that extracts fingerprints from artifact $A$, wherein artifact $A$ captures specific traces during the model training process, independent of content $C$ and style $S$. To further enrich the fingerprint representation, causal fingerprints are defined by weighting and concatenating the differences in projections across multiple semantically invariant embedding spaces, as shown in Equation 2, where $\mathcal{E}$ denotes the set of semantically invariant embedding spaces, $\phi_{s}(\cdot)$ represents the feature extraction function for the $s$th embedding space, $w_{s}$ denotes the weighting coefficient, and $r_{dire}$ denotes the residual computed via the pre-trained Diffusion-based Residual Model (DIRE), detailed in Sec. 2.2.
@@ 2.2 Decouple CF in Semantic-Invariant Latent Space
where $\hat{X}_{dire}$ represents the image reconstructed by the DIRE model, which is used for calculating the causal fingerprint $F_{G}$ in Sec. 2.1.
@@ 2.4 Network Architecture
Here, $x\in\mathbb{R}^{N}$ indicates the original 128-vector feature, while $\hat{x}\in\mathbb{R}^{N}$ denotes the fingerprint feature vector reconstructed by the VAE. $x_{i}$ and $\hat{x}_{i}$ respectively denote the $i$-th element of the original fingerprint feature and the reconstructed fingerprint feature. Additionally, $\mu\in\mathbb{R}^{K}$ denotes the mean vector in the latent space, $\sigma^{2}\in\mathbb{R}^{K}$ denotes the variance vector in the latent space, and the $KL$ divergence between the normally distributed output of the encoder $\mathcal{N}(\mu,\sigma^{2})$ and the standard normal distribution $\mathcal{N}(0,1)$.
Model Attribution and Anonymization. Our attribution network aims to predict the source model of AI-generated images. It takes the causal fingerprint of an image obtained through VAE optimization as input and predicts the identity of its source model. The attribution network we employ utilizes a pre-trained CLIP [21] (ViT-B/32 architecture) enhanced with an attention mechanism [8] to bolster classification capabilities. It is fine-tuned using standard cross-entropy loss. The architecture diagram in Fig. 2 illustrates our model’s artifact extraction, causal fingerprint optimization, visual disentanglement, and attribution process.
In the counterfactual approach, we employ the Projected Gradient Descent (PGD) adversarial perturbation algorithm[17]. By applying perturbations constrained by the causal fingerprint to images, we alter model predictions while maintaining perturbations within the $\mathcal{L}^{\infty}$ norm constraint, achieving anonymization. Its significance lies in: concealing model identity to protect privacy, enhancing attribution accuracy and robustness, isolating causal effects to reduce bias, and supporting reliable generative model development.
@@ 3.1 Experimental Setup
Datasets. To evaluate the performance of model attribution using fingerprints across diverse environments, we constructed the GM-GenImage dataset based on GenImage[32], a large AI-generated image dataset that offers strong content and semantic diversity. We selected 10K images from each of four generative models (BigGAN, ProGAN, Glide, and Stable Diffusion (SD)), yielding a total of 40K images to construct GM-GenImage.
Baseline. We evaluated four major categories of model attribution methods based on pixel-level color, frequency domain features, supervised learning, and manifold learning respectively. We compare representative methods from each group with our approach, including methods based on pixel-level colour: McCloskey et al. [19], and Hae et al. [25]; frequency domain feature-based methods: Dzanic[5]; supervised learning-based methods: Yu et al. [29], and Wang et al. [27]; Euclidean manifold learning-based methods: Hae et al. [24]. Their respective embedding spaces are utilized for comparative analysis.
@@ 3.2 Results
Table 1 presents the model attribution results for fingerprint recognition methods. The convolutional neural network (CNN) classifier outperforms KNN and support vector machines (SVM). Ablation experiments indicate that both the pre-trained ViT model’s class-labeled features and the grayscale FFT low-frequency spatial extraction features lead to poor fingerprint attribution performance. Subsequent multi-space fusion techniques improved attribution performance by reducing weights. Additionally, the FDR metric validated the separability of fingerprint feature spaces, with higher FDR values indicating stronger attribution capability. This method achieves significantly better FDR values than competitors, with an average attribution accuracy improvement of 22.78% (compared to 33.61%, 16.17%, 14%, 15.91%, 23.54%, and 33.42% for other methods), confirming the effectiveness of causal fingerprints in generative model attribution.
@@
Yu et al. (ICCV 2019) [29] 64.43 (+33.61) 148.16 0.657 0.644
Wang et al. (CVPR 2019) [27] 81.87 (+16.17) 3.49 0.837 0.819
McCloskey et al. (arXiv 2018) [19] 84.04 (+14.00) 3.38 0.843 0.840
Dzanic et al. (NeurIPS 2020) [5] 82.13 (+15.91) 48.40 0.824 0.821
Hae et al. (CVPR 2024) [25] 64.62 (+33.42) 1.24 0.699 0.646
Hae et al. (arXiv 2025) [24] 74.5 (+23.54) 83.34 0.858 0.745
@@
multi_fingerprint($F_{0}$) 88.83 199.67 0.888 0.888
multi_fingerprint($F_{1}$) 94.33 312.77 0.944 0.943
multi_fingerprint($F_{2}$) 98.04 357.01 0.980 0.980
@@ 4 Conclusion
5 Acknowledgment
This work was supported by the National Natural Science Foundation of China under Grant No. 62402009, the Key Laboratory of Computing Power Network and Information Security, Ministry of Education under Grant No.2024PY014, and the Science and Technology Development Fund of Macao under Grant No. 0013-2024-ITP1 and 0069/2025/ITP2.